OpenStudy (anonymous):

Finding the sum of the series (Will post the problem below) (I thank you all in advance for contributing)

OpenStudy (anonymous):

\[\sum_{3}^{\inf} \frac{ 1 }{ k(k+4) }\]

OpenStudy (anonymous):

Im not quite understanding how to start a problem. Would I have to separate this by partial sums?

OpenStudy (anonymous):

most likely it's a telescope series. So yes

OpenStudy (anonymous):

Ok how do I do that?

OpenStudy (anonymous):

Im stuck on that part

OpenStudy (kainui):

Have you already made it into partial fractions, show what you've gotten so far.

OpenStudy (kainui):

After you have the partial fractions, start writing out the first few sums of the series. You'll notice after a few terms that you will see the same terms repeat, only negative.

OpenStudy (anonymous):

Is this correct? \[(\frac{ 1 }{ k }-\frac{ 1 }{ k+4 })\]

OpenStudy (anonymous):

that multiply by 1/4

OpenStudy (kainui):

Almost, except that they're both divided by 4.

OpenStudy (anonymous):

So \[(\frac{ 1 }{ 4k }-\frac{ 1 }{ 4k+16 })\]

OpenStudy (kainui):

Sure, and I'd probably just leave it outside rather than distributing it out like that to make it easier. Just multiply your final answer by 1/4.

OpenStudy (anonymous):

yes, to see the pattern easier, i would suggest take out 1/4 and see the patter of 1/k - 1/(k+4)

OpenStudy (anonymous):

Oh ok

OpenStudy (kainui):

\[\frac{1}{4}(\sum_{k=3}^{\infty} \frac{1}{k}-\sum_{k=3}^{\infty} \frac{1}{k+4})\] So here it is easier to see split up. Notice that after 4 terms with the last sum, you'll be adding the same terms that were just subtracted 4 terms ago by the other term. So you can see how it simplifies to \[\frac{1}{4}\sum_{k=3}^{6} \frac{1}{k}\] Which isn't an infinite sum and easy to add up on your own.

OpenStudy (kainui):

Actually writing out the terms rather than doing it all in summation is probably better for you to do so you understand what I'm saying. Write out the first 4 terms and from then on you'll see that the rest of the terms exactly cancel. =) Try it out and we'll help you finish it.

OpenStudy (anonymous):

ok so the first 4 terms of the partial sum are (4/21)+(1/8)+(4/45)+(1/15)

OpenStudy (anonymous):

When they are added up